Circular No. 6766

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997do IN UGC 3845 Y. L. Qiu, Q. Y. Qiao, W. D. Li, and J. Y. Hu, Beijing Astronomical

Observatory (BAO); and A. Esamdin, Urumqi Astronomical Station, on behalf of the BAO supernova survey, report their discovery of a supernova found on unfiltered CCD images with the BAO 0.6-m reflector on Oct. 31 UT. Measured R magnitudes: Oct. 31.84, 15.8; Nov. 1.67, 15.6. The supernova is located at α = 7^{h}26^{m}42^{s}.50, δ = +47^{o}05′36′′.0 (equinox 2000.0), which is 2′′.6 west and 3′′.8 south of the center of UGC 3845. Unfiltered CCD images of the same field taken on Oct. 25 (limiting mag ∼ 18.5) show no star at the position of the supernova. A low-dispersion spectrum, taken on Nov. 1.80 by Qiu, Esamdin, and K. L. Huang (Nanjing Normal University) with the 2.16-m telescope at Xinglong Station, shows this to be a type-Ia supernova before maximum. The spectrum displays a strong Si II 615nm absorption feature. The spectrum is similar to that of SN 1990N, one week before maximum. The recession velocity, derived from the narrow Hα emission line, is 3240 km/s.

AO 0235+164 J. Webb and E. Howard, Florida International University and the

SARA Observatory; and T. Balonek, Foggy Bottom Observatory (FBO), Colgate University, report: “This BL Lac object is currently undergoing a period of optical activity. CCD observations with the SARA 0.9-m telescope at Kitt Peak and the FBO 0.4-m telescope indicate that it has brightened by 0.9 mag in four days. Preliminary R magnitudes documenting its activity are: Oct. 2.24 UT, 17.34 ± 0.07; 12.27, 17.67 ± 0.10; 20.29, 17.16 ± 0.07; 21.29, 17.51 ± 0.04; 23.31, 17.12 ± 0.05; 25.24, 16.62 ± 0.04; 26.17, 16.56 ± 0.03; 29.12, 16.45 ± 0.05; 30.28, 15.99 ± 0.10. The source was typically fainter than mag 18.0 during the past few years. Previous outbursts have exhibited large-amplitude rapid flaring at optical wavelengths, and possible correlated activity with other wavelengths. Further observations of this blazar are encouraged at all wavelengths.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: Sept. 28.83 UT, 11.6 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); Oct. 6.78, 11.9 (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector); 11.88, 11.5 (J. M. Trigo, Benicòssim, Castellon, Spain, 0.17-m reflector); 20.82, 11.2 (G. Comello, Roden, The Netherlands, 0.30-m reflector); 31.77, 11.5 (Bouma).

1997 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green